# mode=buildnew

### 20 Mar 2024 v1

@@ Abstract
When a neural network parameterized loss function consists of many terms, the combinatorial choice of weight multipliers during the optimization process forms a challenging problem. To address this, we proposed a probabilistic graphical model (PGM) for the joint model parameter and multiplier evolution process, with a hypervolume based likelihood that promotes multi-objective descent of each loss term. The corresponding parameter and multiplier estimation as a sequential decision process is then cast into an optimal control problem, where the multi-objective descent goal is dispatched hierarchically into a series of constraint optimization sub-problems. The subproblem constraint automatically adapts itself according to Pareto dominance and serves as the setpoint for the low level multiplier controller to schedule loss landscapes via output feedback of each loss term. Our method is multiplier-free and operates at the timescale of epochs, thus saves tremendous computational resources compared to full training cycle multiplier tuning. We applied it to domain invariant variational auto-encoding with 6 loss terms on the PACS domain generalization task, and observed robust performance across a range of controller hyperparameters, as well as different multiplier initial conditions, outperforming other multiplier scheduling methods. We offered modular implementation of our method, admitting custom definition of many loss terms for applying our multi-objective hierarchical output feedback training scheme to other deep learning fields.
@@ 1 Introduction
In many deep learning fields including domain generalization (Gulrajani and Lopez-Paz, 2020), the loss function for neural networks amounts to a summation of multiple terms, with a multiplier weighting each term. As the number of terms increases, the space of choice for those multipliers grows dramatically. In addition, when the validation set is drawn from a different distribution than the target domain, the validation performance does not faithfully reflect the target domain performance and can easily reach and maintain a saturated value during training. All these factors complicate and add difficulties to the process of model selection using conventional hyperparameter tuning (Probst et al., 2019) methods.
Low level optimization algorithms for descending single objective neural network losses (e.g. a weighted summation of many terms), such as ADAM (Kingma and Ba, 2015), have been widely used, although their theoretical behavior was analyzed predominantly with convex optimization theory (Kingma and Ba, 2015; Reddi et al., 2019), which does not hold globally (Milne, 2019). From a control theory point of view, these low level optimization algorithms drive the model parameters iteratively, which forms a dynamic system. Due to lack of comprehensive theoretical underpinnings of the loss function property with respect to model parameters, this dynamic system has unreliable open-loop behavior. When considering multi-objective optimization of the many loss terms, this lack of theoretical understanding becomes more pronounced.
Inspired by feedback control theory (Doyle et al., 2013) and feedback optimization in power systems (Picallo et al., 2020; He et al., 2023), we treat the dynamic system of model parameters driven by the low level optimization algorithm as an uncontrolled plant, with model parameter as state and the many loss terms as output. We use state dependent multiplier induced loss landscapes as control input and designed a hierarchical control structure (Dietterich et al., 1998; Kulkarni et al., 2016; Sun et al., 2020) to drive the system towards a multi-objective descent goal, without modifying the internal dynamic of low level neural network optimization algorithms.
@@
  • We proposed a probabilistic graphical model (PGM) for depicting the joint multiplier and model parameter adaptation process with a likelihood function imposed, promoting multi-objective descent of each loss term in structured risk. We proposed to transform the problem of parameter inference of this PGM into a sequential decision process through an optimal control formulation.
  • We proposed to solve the sequential decision process for the multi-objective descent goal via breaking the goal into a series of constraint optimization sub-goals which forms a hierarchical controller that automatically adjusts the value of both the multiplier and the model parameters during the training process. The constraint bound updates itself via Pareto dominance. The multi-objective descent of the empirical risk and multidimensional penalty (regularization) loss in Equation 1 offers additional model selection criteria along with validation set performance. Compared to Bayesian optimization, our controller operates at the timescale of model parameter dynamic epoch instead of a complete training, which is more computationally efficient.
@@ 2.4 Domain generalization
2.5 Related work
Multipliers warmup (Sønderby et al., 2016; Ilse et al., 2020) or phase-in (Sicilia et al., 2023) requires ultimate multiplier values for each loss term, thus necessitates hyperparameter (ultimate multiplier values) optimization. Other works, such as (Rezende and Viola, 2018; Klushyn et al., 2019), has introduced constraint optimization by representing the weights of loss terms using Lagrange multipliers, designed to prevent over-regularisation of a single term within the loss function, which is extended to multiple loss terms in (Chen et al., 2022). In addition, Shao et al. (2020) (only deals with 2 loss terms),Chen et al. (2022) proposed similar multiplier adaptation schemes as in Section 3.2.3 where the multiplier changes with a rate proportional to how the corresponding loss component distances from a constraint bound. However, their methods relied on a fixed constraint bound  for each loss term and fixed proportionate gain for each loss term (defined to be $K_{I}$ in Equation 48 in our work), necessitating a hyperparameter search, which is challenging when the number of terms in the loss increases. Moreover, for some regularizers, such as the ones based on KL divergence, it is challenging to estimate the constraint bound range. Specifically in the field of domain generalization, it is not obvious how to choose the constraint bound value and other hyperparameters due to the lack of observations from the target domain, where the validation set from the training domain can easily reach a saturated value due to the high expressive power of modern neural networks.
@@ 3.1 Probabilistic graphical model for parameter and multiplier adaptive optimization process
We propose Figure 2 to describe the sequential decision process of model parameter and multiplier adaptation in the form of a probabilistic graphical model (PGM) (Koller and Friedman, 2009). In the upper part of Figure 2, we model the generative process of $N$ observations of the instance $X$ (e.g. an image) and the corresponding supervision signal $Y$ (e.g. a class label). The path $\beta\rightarrow D_{ob}\rightarrow X$, where, $D_{ob}$ is the data-site Sun et al. (2019a) specific prior parameter (domain variable, characterizing distribution) with hyper-prior $\beta$, models the information which is not captured by the supervision signal $Y$ of $X$.
@@ 4.1 Experimental Setting
This section presents our experiments training the domain generalization model DIVA (Ilse et al., 2020) with $6$ loss terms defined in Equation 4 on the widely used domain generalization benchmark dataset PACS (Li et al., 2017) with different training schemes. Given that the sketch domain within the PACS dataset is considered the most inherently challenging, it was chosen as the sole leave-one-out domain to test the performance of domain generalization. We run experiments to answer the following questions:
@@ 5 Conclusion
This work addresses the issue of combinatorial choice for multidimensional multipliers weighting many terms in neural network loss function. We proposed a probabilistic graphical model for joint multiplier and model parameter optimization with respect to a multi-objective descent of all loss terms. The estimation of model parameter leads to an automatic adjustment scheme for the multipliers, adopting a hierarchical control scheme, which breaks the multi-objective optimization problem into a series of constraint optimization sub-problems. Each sub-problem is configured with a self-adaptive multi-objective setpoint updated via Pareto dominance. A PI-like multiplier controller drives the loss term to satisfy the setpoint constraint for each sub-problem. Our method operates at the timescale of epoch level during training, thus circumvents the need for exhaustive multiplier search and saves tremendous computational resources compared to methods like Bayesian optimization. Our method demonstrates robust generalization performance in the PACS domain generalization task with Domain Invariant Variational Autoencoding Ilse et al. (2020), compared to other multiplier choice or scheduling schemes which produce very unstable behavior in the training due to the need for a combinatorial choice of multipliers. Our method is multiplier-free since the multiplier is the output of our controller, and it has robust performance with respect to different multiplier initial values. For future work, we are interested in testing the performance of our hierarchical feedback optimization on other domain generalization methods, as well as on other deep learning fields in the form of Equation 1. Theoretically, it is interesting to validate Conjecture 3.1, derive sufficient conditions, incorporate more mathematical analysis into multidimensional case for $R$, and investigate the potential $\mu$ convergence behavior.
@@ 6 Contributions And Acknowledgement
XS proposed the idea, developed the theories and algorithms, implemented the code, designed and carried out the experiments, processed experimental data and generated the figures, wrote the manuscript. NC proposed the PI-like multiplier adaptation in Section 3.2.3, helped XS with PI-like controller implementation, held various discussions, proofread the manuscript. AG initiated the visualization code for Figure 5 and Figure 6, refined by XS, proofread and improved the manuscript. YX developed Corollary 3.2 and corresponding remarks with XS, proofread and improved the constraint optimization part of the theory, discussed on control theory part. CF initiated the benchmark code with refinements from XS, proofread Corollary 3.1. ED and FD tested and improved the code for benchmark, proofread the manuscript. FD further improved the tensorboard visualization of the algorithm. LB proofread Corollary 3.1, tested the benchmark code. DS tested the benchmark code, assisted XS in collecting experimental results. CM supervised the project, offered intensive and in depth proofread of the manuscript with major quality improvements.
We sincerely thank Dr. Changxin Liu and Dr. Prof. Jann Rolfes for spending hours on proofreading the manuscript and offering many helpful suggestions to the manuscript. We thank Tomas Chobola for Inkscape support.

### 10 Apr 2024 v2

@@ Abstract
We address the online combinatorial choice of weight multipliers for multi-objective optimization of many loss terms parameterized by neural works via a probabilistic graphical model (PGM) for the joint model parameter and multiplier evolution process, with a hypervolume based likelihood promoting multi-objective descent. The corresponding parameter and multiplier estimation as a sequential decision process is then cast into an optimal control problem, where the multi-objective descent goal is dispatched hierarchically into a series of constraint optimization sub-problems. The subproblem constraint automatically adapts itself according to Pareto dominance and serves as the setpoint for the low level multiplier controller to schedule loss landscapes via output feedback of each loss term. Our method is multiplier-free and operates at the timescale of epochs, thus saves tremendous computational resources compared to full training cycle multiplier tuning. It also circumvents the excessive memory requirements and heavy computational burden of existing multi-objective deep learning methods. We applied it to domain invariant variational auto-encoding with 6 loss terms on the PACS domain generalization task, and observed robust performance across a range of controller hyperparameters, as well as different multiplier initial conditions, outperforming other multiplier scheduling methods. We offered modular implementation of our method, admitting extension to custom definition of many loss terms.
@@ 1 Introduction
In many deep learning fields including domain generalization (Gulrajani and Lopez-Paz, 2020), the loss function for neural networks amounts to a summation of multiple terms, with a multiplier weighting each term. As the number of loss terms increases, the space of choice for those multipliers grows dramatically. In addition, when the validation set is drawn from a different distribution than the target domain, the validation performance does not faithfully reflect the target domain performance (Chen et al., 2022b) and can easily reach and maintain a saturated value during training. All these factors complicate and add difficulties to the process of model selection using conventional hyperparameter tuning (Probst et al., 2019) methods.
Instead of optimizing a fixed weighted summation of many neural network loss terms using a single objective optimization algorithm which relies on heavy hyperparameter tuning of the weight multipliers, multi-objective optimization on these loss terms has raised attention in the deep learning community recently (Sener and Koltun, 2018; Lin et al., 2019; Mahapatra and Rajan, 2021; Chen et al., 2022b). Previous works in this direction (Mahapatra and Rajan, 2021; Chen et al., 2022b) addresses the problem via careful choice of multipliers at each optimization iteration instead of fixed for the whole training cycle, but suffers from excessive memory requirement and heavy extra computation burden, which hinders their applicability to large dataset and large scale neural networks Chen et al. (2022b). Furthermore, it is restricted to Exact Pareto optimization (Mahapatra and Rajan, 2021) where the solution should be aligned to a user input preference vector, while how to set that is not clear in the general case. Their performance further relies on careful tuning of learning rate of single objective low level optimization algorithms Chen et al. (2022b). Single objective optimization algorithms like ADAM (Kingma and Ba, 2015) were analyzed predominantly with convex optimization theory (Kingma and Ba, 2015; Reddi et al., 2019), which does not hold globally (Milne, 2019). From a control theory point of view, these low level optimization algorithms drive the model parameters iteratively, which forms a dynamic system. Due to lack of comprehensive theoretical underpinnings of the loss function property with respect to model parameters, this dynamic system has unreliable open-loop behavior.
Inspired by feedback control theory (Doyle et al., 2013) and feedback optimization in power systems (Picallo et al., 2020; He et al., 2023; Hauswirth et al., 2024), we treat the dynamic system of model parameters driven by the low level optimization algorithm as an uncontrolled plant, with model parameter as state and the many loss terms as output. We use state dependent multiplier induced loss landscapes as control input and designed a hierarchical control structure (Dietterich et al., 1998; Kulkarni et al., 2016; Sun et al., 2020) to drive the system towards a multi-objective descent goal, without modifying the internal dynamic of low level neural network optimization algorithms.
@@
  • We proposed a probabilistic graphical model (PGM) for depicting the joint multiplier and model parameter adaptation process with a likelihood function imposed, promoting multi-objective descent of each loss term, which can be used to described general multi-objective deep learning. We proposed to transform the problem of parameter inference of this PGM into a sequential decision process through an optimal control formulation.
  • We proposed to solve the sequential decision process for the multi-objective descent goal via breaking the goal into a series of constraint optimization sub-goals which forms a hierarchical controller that automatically adjusts the value of both the multiplier and the model parameters during the training process. The constraint bound updates itself via Pareto dominance. The multi-objective descent of the empirical risk and multidimensional penalty (regularization) loss in Equation 1 offers additional model selection criteria along with validation set performance. Compared to Bayesian optimization, our controller operates at the timescale of model parameter dynamic epoch instead of a complete training, which saves tremendous computational resources. Compared to existing multi-objective deep learning methods, it circumvent excessive memory requirements and heavy extra computational burden.
@@ 3.1 Probabilistic graphical model for parameter and multiplier adaptive optimization process
Gradient based optimization algorithms for neural network defines a dynamic system for model parameters where the weight multiplier in Equation 1 configures such a dynamic system. Instead of using fixed multipliers (or fixed multipliers’ ultimate values), we propose to choose the multiplier values also dynamically. Due to the stochastic nature of the gradient based optimization process in deep learning with different subsample and mini-batch distributions for each iteration, it is appropriate to use a Probabilistic Graphical Model (PGM) to describe such a stochastic process. PGM has been used to describe sequential decision processes like reinforcement learning (Levine, 2018; Sun and Bischl, 2019). We propose Figure 2 to describe the sequential decision process of model parameter and multiplier adaptation in the form of a probabilistic graphical model (PGM) (Koller and Friedman, 2009). In the upper part of Figure 2, we model the generative process of $N$ observations of the instance $X$ (e.g. an image) and the corresponding supervision signal $Y$ (e.g. a class label). The path $\beta\rightarrow D_{ob}\rightarrow X$, where, $D_{ob}$ is the data-site Sun et al. (2019a) specific prior parameter (domain variable, characterizing distribution) with hyper-prior $\beta$, models the information which is not captured by the supervision signal $Y$ of $X$.
@@
The PGM we defined here can also be used to describe other multi-objective deep learning methods Mahapatra and Rajan (2021) via online combinatorial choice of multipliers.
@@ 3.2.4 Multi-objective setpoint based model selection
4 Related work and Discussion
Multipliers warmup (Sønderby et al., 2016; Ilse et al., 2020) or phase-in (Sicilia et al., 2023) requires ultimate multiplier values for each loss term, thus necessitates hyperparameter (ultimate multiplier values) optimization. Other works, such as (Rezende and Viola, 2018; Klushyn et al., 2019), has introduced constraint optimization by representing the weights of loss terms using Lagrange multipliers, designed to prevent over-regularisation of a single term within the loss function, which is extended to multiple loss terms in (Chen et al., 2022a). In addition, Shao et al. (2020) (only deals with 2 loss terms),Chen et al. (2022a) proposed similar multiplier adaptation schemes as in Section 3.2.3 where the multiplier changes with a rate proportional to how the corresponding loss component distances from a constraint bound. However, their methods relied on a fixed constraint bound  for each loss term and fixed proportionate gain for each loss term (defined to be $K_{I}$ in Equation 48 in our work), necessitating a hyperparameter search, which is challenging when the number of terms in the loss increases. Moreover, for some regularizers, such as the ones based on KL divergence, it is challenging to estimate the constraint bound range. Specifically in the field of domain generalization, it is not obvious how to choose the constraint bound value and other hyperparameters due to the lack of observations from the target domain, where the validation set from the training domain can easily reach a saturated value due to the high expressive power of modern neural networks.
Earlier multi-objective deep learning (Sener and Koltun, 2018; Lin et al., 2019) suffers from increased computational cost with the number of losses, explicit computation of the gradient with respect to each loss term and gradient norms, making it practically infeasible Mahapatra and Rajan (2021) for large neural networks and large datasets. For instance, Pareto Invariance Risk Minimization (PAIR) (Chen et al., 2022b) studies domain generalization with the use of the multi-objective optimization technique Exact Pareto Optimal Search (EPOS) (Mahapatra and Rajan, 2021), which requires explicit full gradient information for each loss term and computation of the $C$ matrix with complexity $m^{2}n$ (Mahapatra and Rajan, 2021). With $m$ being the number of loss terms (equivalent to $d+1$ in our paper), and $n$ the dimension of gradient, which is super high in modern neural networks (e.g. 200 million for ResNet50), thus results in heavy memory requirements and excessive computation burden, and hinders the deployment to large-scale networks Chen et al. (2022b). In comparison, our multi-objective optimization method (M-HOF-Opt) need only $d$ scalar multiplications, thus the extra computation is negligible and can be used in large-scale neural networks. EPOS (Mahapatra and Rajan, 2021) also requires the user to provide a preference vector which can be difficult to set, say with 6 objectives for our case. In addition, Chen et al. (2022b) reported hyperparameter tuning for the preference vector and step length is needed for the method to work properly. In comparison, our method, M-HOF-Opt is preference vector free. It ensures the Pareto descent due to the update rule of our setpoint. There is no requirement to search for step length.
Note that the EPOS method Mahapatra and Rajan (2021) is searching for Exact Pareto optimal (Mahapatra and Rajan, 2021), which might not exist. Chen et al. (2022b) divides the training into two phases where the first phase uses single-objective optimization favoring ERM loss while the second phase is for balancing different loss terms. However, how many computation resources should one set for the first phase is not clear. Furthermore, the first stage (phase) can affect the exact Pareto optimality.
@@ 5.1 Experimental Setting
This section presents our experiments training the domain generalization model DIVA (Ilse et al., 2020) with $6$ loss terms defined in Equation 4 on the widely used domain generalization benchmark dataset PACS (Li et al., 2017) with different training schemes. Due to the excessive memory requirements and heavy computation of earlier multi-objective deep learning methods discussed in Section 4, which hinders their application, we can only compare our training scheme with other multiplier scheduling techniques. Given that the sketch domain within the PACS dataset is considered the most inherently challenging, it was chosen as the sole leave-one-out domain to test the performance of domain generalization. We run experiments to answer the following questions:
@@ 6 Conclusion
This work addresses the issue of combinatorial choice for multidimensional multipliers weighting many terms in neural network loss function. We proposed a probabilistic graphical model for joint multiplier and model parameter optimization with respect to a multi-objective descent of all loss terms. The estimation of model parameter leads to an automatic adjustment scheme for the multipliers, adopting a hierarchical control scheme, which breaks the multi-objective optimization problem into a series of constraint optimization sub-problems. Each sub-problem is configured with a self-adaptive multi-objective setpoint updated via Pareto dominance. A PI-like multiplier controller drives the loss term to satisfy the setpoint constraint for each sub-problem. Our method operates at the timescale of epoch level during training, thus circumvents the need for exhaustive multiplier search and saves tremendous computational resources compared to methods like Bayesian optimization. It also circumvents the excessive memory requirements and heavy computational burden of existing multi-objective deep learning methods. Our method demonstrates robust generalization performance in the PACS domain generalization task with Domain Invariant Variational Autoencoding Ilse et al. (2020), compared to other multiplier choice or scheduling schemes which produce very unstable behavior in the training due to the need for a combinatorial choice of multipliers. Our method is multiplier-free since the multiplier is the output of our controller, and it has robust performance with respect to different multiplier initial values. For future work, we are interested in testing the performance of our hierarchical feedback optimization on other domain generalization methods, as well as on other deep learning fields in the form of Equation 1. Theoretically, it is interesting to validate Conjecture 3.1, derive sufficient conditions, incorporate more mathematical analysis into multidimensional case for $R$, and investigate the potential $\mu$ convergence behavior.
@@ 7 Contributions And Acknowledgement
XS proposed the idea, developed the theories and algorithms, implemented the code, designed and carried out the experiments, processed experimental data and generated the figures, wrote the manuscript. NC proposed the PI-like multiplier adaptation in Section 3.2.3, helped XS with PI-like controller implementation, held various discussions, proofread and improved the manuscript. AG initiated the visualization code for Figure 5 and Figure 6, refined by XS, proofread and improved the manuscript. YX developed Corollary 3.2 and corresponding remarks with XS, proofread and improved the constraint optimization part of the theory, discussed on control theory part. CF initiated the benchmark code with refinements from XS, proofread Corollary 3.1. ED and FD tested and improved the code for benchmark, proofread the manuscript. FD further improved the tensorboard visualization of the algorithm. LB proofread Corollary 3.1, tested the benchmark code. DS tested the benchmark code, assisted XS in collecting experimental results. CM supervised the project, offered intensive and in depth proofreads of the manuscript with major quality improvements.
We sincerely thank Dr. Changxin Liu and Dr. Prof. Jann Rolfes for spending hours on proofreading the manuscript and offering many helpful suggestions to the manuscript. We thank Tomas Chobola for Inkscape support. We thank the anonymous reviewers for offering suggestions to enhance the readability of our paper.